OpenStudy (anonymous):

y=1-3x graph the question

OpenStudy (anonymous):

start by find the x and y intercepts

OpenStudy (anonymous):

ya but grpah it i dont know

OpenStudy (anonymous):

hurry

OpenStudy (karatechopper):

let x be 0 for y int y=1-3(0) y=1-0 y=1 (0,1) let y be 0 for x int 0=1-3x 0-1=1-1-3x -1=-3x -1/-3=-3x/-3 1/3=x (1/3, 0) your two points will be (0,1) and (1/3,0) now for the graph......i will put tht on here hold on

OpenStudy (anonymous):